OpenStudy (anonymous):

Find the zeros of the function by factoring. f(x)=9x^4-25x^2

OpenStudy (anonymous):

\[(3x^2)^2-(5x)^2\] \[(3x^2+5x)(3x^2-5x)\] since \[a^2-b^2 = (a+b)(a-b)\]

OpenStudy (anonymous):

got it phuelnora

OpenStudy (anonymous):

uhh not reallly

OpenStudy (anonymous):

what are the zeros?

OpenStudy (anonymous):

ok now see here : \[3x^2+5x=0\] \[x(3x+5)=0\] \[x=o , x=-5/3\] \[3x^2-5x=0\] \[x(3x-5)=0\] \[x=0 , x=5/3\] zeros are 0 , 5/3 and -5/3

OpenStudy (anonymous):

now got it

OpenStudy (anonymous):

thank you